OpenStudy (anonymous):

Solve The Equation 1/3 (6+9x)

TheSmartOne (thesmartone):

\[\huge \color{blue}{Welcome~To~OpenStudy!!!}\]

TheSmartOne (thesmartone):

It looks like you are new here. This tutorial can help you get started. http://openstudy.com/study#/updates/543de42fe4b0b3c6e146b5e8

TheSmartOne (thesmartone):

\[\frac{ 6+9x }{ 3 }=2+3x\]